OpenStudy (anonymous):

if log (a+b/2)=1/2 (log a+log b) prove that a = b.

OpenStudy (anonymous):

IF \[\log{(a+b/2)}=1/2 (\log{a}+\log{b})\] PROVE THAT A=B

OpenStudy (asnaseer):

firstly, how can you combine:\[\log(a)+\log(b)\]using the standard log rules?

OpenStudy (anonymous):

LOG AB

OpenStudy (asnaseer):

correct. next how can we take the 1/2 from outside the log to the inside, i.e:\[\frac{1}{2}\log(ab)\]using standard log rules?

OpenStudy (anonymous):

\[\log{ab^1/2} \]

OpenStudy (asnaseer):

to be clearer, its actually:\[\log(ab)^\frac{1}{2}\]

OpenStudy (anonymous):

log ab^1/2 or\[\log{\sqrt{ab}} \]

OpenStudy (asnaseer):

correct

OpenStudy (asnaseer):

so now we are left with:\[\log(\frac{a+b}{2})=\log(\sqrt{ab})\]

OpenStudy (asnaseer):

which means we can equate the things inside the logs

OpenStudy (anonymous):

yes i tried but it didn't work out

OpenStudy (asnaseer):

we get:\[\frac{a+b}{2}=\sqrt{ab}\]correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now?

OpenStudy (asnaseer):

so now square both sides

OpenStudy (anonymous):

i did that it didn't work out could u tell

OpenStudy (asnaseer):

what do you get when you square both sides?

OpenStudy (anonymous):

\[(a^2+2ab+b^2)/4=ab\]

OpenStudy (asnaseer):

now multiply both sides by 4 - what do you get?

OpenStudy (anonymous):

\[a^2+2ab+b^2=4ab\]

OpenStudy (asnaseer):

correct - almost there now. now subtract 4ab from both sides and what do you get?

OpenStudy (anonymous):

\[a^2-2ab+b^2=0\]

OpenStudy (asnaseer):

perfect! now can you see how to factorise \(a^2-2ab+b^2\)?

OpenStudy (anonymous):

\[a^2-b^2=0\]

OpenStudy (asnaseer):

not quite, it should be:\[(a-b)^2=0\]

OpenStudy (anonymous):

\[a^2=b^2\] so a=b

OpenStudy (asnaseer):

your factorisation was incorrect

OpenStudy (anonymous):

oh sorry

OpenStudy (anonymous):

(a-b)^2

OpenStudy (asnaseer):

yup

OpenStudy (anonymous):

now?

OpenStudy (asnaseer):

so you end up with:\[(a-b)^2=0\]which means?

OpenStudy (asnaseer):

if "something" squared is zero, then that "something" must be?

OpenStudy (anonymous):

0

OpenStudy (asnaseer):

correct, so we get:\[(a-b)^2=0\implies a-b=0\]

OpenStudy (anonymous):

since a-b=0 a=b

OpenStudy (asnaseer):

thats it - you have proved it!

OpenStudy (anonymous):

actually i wrote it earlier but forgot to post :p

OpenStudy (asnaseer):

:)